The volume of a sphere is numerically equal to the surface area of the sphere. If this value is equal to Nπ, what is the value of N?
Answer by mananth(16946)

vol of sphere = 4/3 pir^3
surface area = 4 pir^2
.
.
.
4/3 pi r^3= N pi
N= 4/3 r^3
.
.

N pi = 4 pi r^2
N= 4r^2
.
.
equate both
4r^2=4/3 r^3
r=3
N= 4r^2
N= 4*3*3
N=36